OpenStudy (yourgirlbri):

Situation: A student in Greece discovers a pottery bowl that contains 65% of its original amount of C-14. N0=Initial amount of C-14 (at time t=0) N= amount of C-14 at time t t= time, in years k=0.0001 Find the age of the pottery bowl to the nearest year.

OpenStudy (abhisar):

Hi @yourgirlbri !

OpenStudy (abhisar):

You can use the following formula \(\sf N=N_0e^{-kt}\)

OpenStudy (yourgirlbri):

I don't understand how to plug it in :(

OpenStudy (abhisar):

Are you familiar with logarithms and natural logarithms?

OpenStudy (yourgirlbri):

not really.. I have a calculator though and I've just been plugging stuff in

OpenStudy (abhisar):

Fine, I'll show you how to plug in values for this question.

OpenStudy (yourgirlbri):

7340? (somebody told me this was the best answer idk)

OpenStudy (abhisar):

Given, N0=Initial amount of C-14 (at time t=0) N= amount of C-14 at time t k=0.0001 To find, t= time, in years \(\sf N=N_0e^{-kt}\\ \Rightarrow N/N_0=e^{-kt} \\ \Rightarrow 0.65=e^{-0.0001t}\) Now taking natural logs both sides we get ln(0.65)=-0.0001 x t Solve for t now

OpenStudy (yourgirlbri):

I literally don't understand at all... and I have 20 questions due tomorrow fml

OpenStudy (abhisar):

What part you don't understand?

OpenStudy (yourgirlbri):

honestly, all of it

OpenStudy (abhisar):

You simply have to plug in the given values to find the unknown one. Try reading what I posted again.

OpenStudy (yourgirlbri):

are you multiplying k and t?

OpenStudy (yourgirlbri):

I got 4308 is that right?

OpenStudy (abhisar):

Yes, It's right!

OpenStudy (yourgirlbri):

is there an easier way to easily plug in with other questions similar?

OpenStudy (abhisar):

I am afraid there is not in my knowledge yet.